OpenStudy (anonymous):

write an equation of a line in point slope form that has a slope of -2 and passes through (5,-1)

OpenStudy (ajprincess):

for a line passin through a point (x1,y1) and has a slope of m the equation will be \[\frac{y-y_1}{x-x_1}=m\] \[y-y_1=m(x-x_1)\] \[y=mx-mx_1+y_1\] Nw can u do?